LIMITING REAGENT & % Yield Problem Set

At high temperatures, sulfur combines with iron to form the brown-black iron (II) sulfide:

Fe (s) + S (l) → FeS (s)

In one experiment, 7.62 g of Fe are allowed to react with 8.67 g of S.

a.  What is the limiting reagent, and what is the reactant in excess         

b.  Calculate the mass of FeS formed.   


1
Expert's answer
2021-04-21T05:59:59-0400

Findtheamountofsubstanceofbothreagents.n(Fe)=m/Ar=7.62/56=0.14moln(S)=8.67/32=0.27mola.Sisreagentinexcessb.m(FeS)=nMr=0.14∗88=12.32g